OpenStudy (anonymous):

graph the system of equations and state the solution: -3x+y=-3 y=x-3

OpenStudy (anonymous):

there are 2 equations: -3x+y=-3 and y=x-3

OpenStudy (zzr0ck3r):

did you graph them?

OpenStudy (anonymous):

well, i got the first equation to look like: y=3x-3. i haven't graphed them yet. how do i do that?

OpenStudy (anonymous):

i also have to state the solution

OpenStudy (zzr0ck3r):

wo we knwo y=3x-3 has a slope of 3 and a y intercept at y = -3 so start on the point where y = -3 and go up 3 and over 1

OpenStudy (anonymous):

ok...

OpenStudy (zzr0ck3r):



OpenStudy (anonymous):

for x, i got (1,0)

OpenStudy (zzr0ck3r):

and for the next one we know the slope is 1 and y intercept at -3 so start and -3 and go up 1 and over 1

OpenStudy (zzr0ck3r):

well and we know the solution is (0,-3) because these are linear and they will only have one solution. and we know they both have the same y intercept so the y intercept is the solution

OpenStudy (anonymous):

so, y is (0,-3) and x is (1,-2)

OpenStudy (zzr0ck3r):

we can check this with algebra y = x-3 y = 3x-3 so 3x-3 = x-3 so 3x = x so 3x-x=0 so x(3-1) = 0 so x(2) = 0 and thus x = 0 when x = 0 both 3x-3 and x-3 are -3 thus a slution at (0,-3)

OpenStudy (zzr0ck3r):

the solution is the pair (x,y) is the point when the two lines cross, so x = 0, and y = -3 so (0,-3)

OpenStudy (anonymous):

so, the solution is just (0,-3), right?

OpenStudy (zzr0ck3r):